OpenStudy (anonymous):

16z^2/2z^2-8z+8*2z-4/4z

OpenStudy (anonymous):

8z^4+7z

OpenStudy (anonymous):

far from what I got! what did you do? please

jimthompson5910 (jim_thompson5910):

16z^2 = (4z)(4z), so one 4z up top cancels with the 4z down below 2z^2-8z+8 = 2(z-2)(z-2) and 2z-4 = 2(z-2) So the 2's outside the parenthesis cancel and one copy of 'z-2' cancels This leaves you with \[\large \frac{4z}{z-2}\]

OpenStudy (anonymous):

i must have copied incorrectly

jimthompson5910 (jim_thompson5910):

it's \[\large \frac{16z^2}{2z^2-8z+8}\times\frac{2z-4}{4z}\]

OpenStudy (anonymous):

oooooooooooh ok sorry

jimthompson5910 (jim_thompson5910):

yeah I had to look the problem 100 times over to make sure lol

OpenStudy (anonymous):

thanks! ;)